# When isometry and equivalence for skew constacyclic codes coincide

### v1

@@ 6. On non monomial homomorphisms
Having established the scarcity of monomial homomorphisms, we now show further that if $S$ is an integral domain (e.g., a field), then in a large number of cases every homomorphism between proper nonassociative algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is Hamming weight preserving — in fact, monomial of degree one, that is, is of the form $G_{\tau,\alpha}$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$ and nonzero $\alpha\in S$.
@@
Suppose $f(t)\in S[t;\sigma]$ is a monic polynomial of degree $\mathbf{n}$ and $b\in S^{\times}$. Suppose that $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a nonzero homomorphism such that $G(t)$ is not a monomial, whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$. Then $\mathbf{m}|\mathbf{n}$ and $G(t)=p(t^{\mathbf{m}})t$ for some $p\in S[x]$ of degree less than $\frac{\mathbf{n}}{\mathbf{m}}$.
@@
for some $c_{i}\in S$. As in the proof of Lemma 4.3, which holds for any commutative unital ring $\S$, the condition $G(td)=G(\sigma(d)t)$ for all $d\in S$ implies that $c_{i}=0$ for all $i$ such that $i\not\equiv 1\mod\mathbf{m}$. Therefore $G(t)$ takes the form given in the statement. It remains to prove that $\mathbf{m}|\mathbf{n}$.
As compared with the monomial case, the proof now is more intricate because each term of $G(t)^{s}$ is a complex polynomial. In the following, we differentiate between the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and the value of $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, that is, after reduction ${\rm mod}_{r}(t^{\mathbf{n}}-b)$ (whose well-definedness is the key question).
Let $k=\deg(G(t))>1$. Let $s>1$ be the least value for which $sk\geq\mathbf{n}$; since $k>1$ we have $s<\mathbf{n}$. Then since $G$ is homomorphism, $G(t)=tp(t^{\mathbf{m}})$ must be power associative for all powers less than $\mathbf{n}$ (since this is true in $S[t;\sigma]/S[t;\sigma]f(t)$) and we have $G(t^{s})=G(t)^{s}$. Furthermore, since $G(t)^{s-1}=(tp(t^{\mathbf{m}}))^{s-1}$ has degree less than $\mathbf{n}$, all monomials appearing with nonzero coefficient in $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ have degree less than $2\mathbf{n}$. Thus we may compute $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ by first evaluating the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and then reducing modulo the relation $t^{\mathbf{n}+r}\equiv t^{r}b\equiv\sigma^{r}(b)t^{r}$.
Write $\mathcal{D}^{s}$ for the set of all $d\in\mathbb{N}$ that occur as the degree of some monomial appearing in the expansion of $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ and $\mathcal{B}^{s}$ for the set of all degrees appearing in $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. Then $\mathcal{B}^{s}\subset\{0,1,\cdots,\mathbf{n}-1\}$ and $\mathcal{D}^{s}\subset\mathcal{B}^{s}+(\mathbf{n}+\mathcal{B}^{s})$.
Since each monomial of $G(t)=tp(t^{\mathbf{m}})$ has degree of the form $1+\ell\mathbf{m}$ for some $\ell\geq 0$, it follows that for each $d\in\mathcal{D}^{s}$ we have $d\equiv s\mod\mathbf{m}$. For each such $d$ that satisfies $d\geq\mathbf{n}$, we obtain a monomial term $\lambda t^{d-\mathbf{n}}$ in the expression for $G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, for some $\lambda\in S$.
@@
If $a,b\in S^{\times}$ and $\mathbf{m}\nmid\mathbf{n}$, then any nonzero homomorphism $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial of degree one.
@@
The theorem shows that there are no polynomial homomorphisms in this case and Corollary 4.5 proves that under the same hypothesis plus any Hamming-weight-preserving homomorphism has degree one. ∎
@@
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial homomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, as in Theorem 6.1. Let $k=\deg(G(t))$. Suppose further that
@@
Then $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ must be associative, that is, $b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
@@
where this is well-defined in $S[t;\sigma]$. If $d<\mathbf{n}$, then $c_{\gamma}$ is also the coefficient corresponding to this monomial product in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, since products with total degree less than $\mathbf{n}$ are associative. If $d>\mathbf{n}$, however, then the value of the coefficient will a priori depend on the order of associators.
By ($\star$6.3), we have $(s+1)k<2\mathbf{n}$. Then $d^{\prime}=d_{max}-\mathbf{n}=(s+1)k-\mathbf{n}$ is the degree of a monomial appearing in $G(t)^{s+1}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ and we wish to compare its coefficients as computed in the two distinct ways ($G(t)G(t)^{s}$ and $G(t)^{s}G(t)$). The coefficient of $t^{d^{\prime}}$ in $G(t)^{s+1}$ in either case is the sum of all monomials in $I_{d^{\prime}}$, plus additional terms that correspond to products in $I_{d_{max}}$ that were reduced by the relation in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to terms of degree $d^{\prime}$. By construction, however, $I_{d_{max}}$ is a singleton, being composed only of the products of the top degree terms:
@@
We evaluate the corresponding term in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ in two ways as follows. First note that
@@
where we have used right division by $t^{\mathbf{n}}-b$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ and that $sk-\mathbf{n}\equiv s\mod\mathbf{m}$. Thus the monomial corresponding to $\gamma_{max}$ in the product $G(t)^{s}G(t)$ is
@@
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, and whose degree $k$ satisfies ($\star$6.3), as in Theorem 6.1. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
@@
Suppose $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$, that is, $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ must be associative.
As mentioned, having such a complete understanding of all homomorphisms between Petit algebras has significant impact to the classification of various associated error-correcting codes. In addition to the conjecture above, we identify that it is possible to extend the results of our paper to the case of noncommutative rings. It would also be interesting to see how our methods can be applied when we look at homomorphisms between general Petit algebras.
Acknowledgments
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation.

### v3

@@ 6. On non monomial homomorphisms
Having established the scarcity of monomial homomorphisms, we now show further that if $S$ is an integral domain (e.g., a field), then in a large number of cases every homomorphism between proper nonassociative rings of the form $\mathbb{S}_{t^{\mathbf{n}}-a}$ is Hamming-weight preserving — in fact, monomial of degree one, that is, is of the form $G_{\tau,\alpha}$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$ and nonzero $\alpha\in S$.
@@
Suppose $f(t)\in S[t;\sigma]$ is a monic polynomial of degree $\mathbf{n}$ and $b\in S^{\times}$. Suppose that $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is a nonzero ring homomorphism whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, such that $G(t)$ is not a monomial. Then $\mathbf{m}|\mathbf{n}$ and $G(t)=p(t^{\mathbf{m}})t$ for some $p\in S[x]$ of degree less than $\frac{\mathbf{n}}{\mathbf{m}}$.
@@
for some $c_{i}\in S$. As in the proof of Lemma 4.3, which holds for any commutative unital ring $S$, the condition $G(td)=G(\sigma(d)t)$ for all $d\in S$ implies that $c_{i}=0$ for all $i$ such that $i\not\equiv 1\mod\mathbf{m}$. Therefore $G(t)$ takes the form given in the statement. It remains to prove that $\mathbf{m}|\mathbf{n}$.
As compared with the monomial case, the proof now is more intricate because each term of $G(t)^{s}$ is a complex polynomial. In the following, we differentiate between the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and the value of $G(t)^{s}\in\mathbb{S}_{t^{\mathbf{n}}-b}$, that is, after reduction ${\rm mod}_{r}(t^{\mathbf{n}}-b)$ (whose well-definedness is the key question).
Let $k=\deg(G(t))>1$. Let $s>1$ be the least value for which $sk\geq\mathbf{n}$; since $k>1$ we have $s<\mathbf{n}$. Then since $G$ is homomorphism, $G(t)=tp(t^{\mathbf{m}})$ must be power associative for all powers less than $\mathbf{n}$ (since this is true in $\mathbb{S}_{f}$) and we have $G(t^{s})=G(t)^{s}$. Furthermore, since $G(t)^{s-1}=(tp(t^{\mathbf{m}}))^{s-1}$ has degree less than $\mathbf{n}$, all monomials appearing with nonzero coefficient in $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ have degree less than $2\mathbf{n}$. Thus we may compute $G(t)^{s}\in\mathbb{S}_{t^{\mathbf{n}}-b}$ by first evaluating the expression $(tp(t^{\mathbf{m}}))^{s}$ in $S[t;\sigma]$ and then reducing modulo the relation $t^{\mathbf{n}+r}\equiv t^{r}b\equiv\sigma^{r}(b)t^{r}$.
Write $\mathcal{D}^{s}$ for the set of all $d\in\mathbb{N}$ that occur as the degree of some monomial appearing in the expansion of $(tp(t^{\mathbf{m}}))^{s}\in S[t;\sigma]$ and $\mathcal{B}^{s}$ for the set of all degrees appearing in $G(t)^{s}\in\mathbb{S}_{t^{\mathbf{n}}-b}$. Then $\mathcal{B}^{s}\subset\{0,1,\cdots,\mathbf{n}-1\}$ and $\mathcal{D}^{s}\subset\mathcal{B}^{s}+(\mathbf{n}+\mathcal{B}^{s})$.
Since each monomial of $G(t)=tp(t^{\mathbf{m}})$ has degree of the form $1+\ell\mathbf{m}$ for some $\ell\geq 0$, it follows that for each $d\in\mathcal{D}^{s}$ we have $d\equiv s\mod\mathbf{m}$. For each such $d$ that satisfies $d\geq\mathbf{n}$, we obtain a monomial term $\lambda t^{d-\mathbf{n}}$ in the expression for $G(t)^{s}\in\mathbb{S}_{t^{\mathbf{n}}-b}$, for some $\lambda\in S$.
@@
If $a,b\in S^{\times}$ and $\mathbf{m}\nmid\mathbf{n}$, then any nonzero homomorphism $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ whose restriction to $S$ is given by some automorphism $\tau$ commuting with $\sigma$, must be monomial of degree one.
@@
The theorem shows that there are no polynomial homomorphisms in this case and Corollary 4.5 proves that under the same hypothesis any Hamming-weight-preserving homomorphism has degree one. ∎
@@
Suppose $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is a non monomial homomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, as in Theorem 6.1. Let $k=\deg(G(t))$. Suppose further that
@@
Then $\mathbb{S}_{t^{\mathbf{n}}-b}$ must be associative, that is, $b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
@@
where this is well-defined in $S[t;\sigma]$. If $d<\mathbf{n}$, then $c_{\gamma}$ is also the coefficient corresponding to this monomial product in $\mathbb{S}_{t^{\mathbf{n}}-b}$, since products with total degree less than $\mathbf{n}$ are associative. If $d>\mathbf{n}$, however, then the value of the coefficient will a priori depend on the order of associators.
By ($\star$6.3), we have $(s+1)k<2\mathbf{n}$. Then $d^{\prime}=d_{max}-\mathbf{n}=(s+1)k-\mathbf{n}$ is the degree of a monomial appearing in $G(t)^{s+1}\in\mathbb{S}_{t^{\mathbf{n}}-b}$ and we wish to compare its coefficients as computed in the two distinct ways ($G(t)G(t)^{s}$ and $G(t)^{s}G(t)$). The coefficient of $t^{d^{\prime}}$ in $G(t)^{s+1}$ in either case is the sum of all monomials in $I_{d^{\prime}}$, plus additional terms that correspond to products in $I_{d_{max}}$ that were reduced by the relation in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to terms of degree $d^{\prime}$. By construction, however, $I_{d_{max}}$ is a singleton, being composed only of the products of the top degree terms:
@@
We evaluate the corresponding term in $\mathbb{S}_{t^{\mathbf{n}}-b}$ in two ways as follows. First note that
@@
where we have used right division by $t^{\mathbf{n}}-b$ in $\mathbb{S}_{t^{\mathbf{n}}-b}$ and that $sk-\mathbf{n}\equiv s\mod\mathbf{m}$. Thus the monomial corresponding to $\gamma_{max}$ in the product $G(t)^{s}G(t)$ is
@@
Suppose $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$, and whose degree $k$ satisfies ($\star$6.3), as in Theorem 6.1. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$.
@@
Suppose $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is a non monomial isomorphism satisfying $G|_{S}=\tau$ for some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$. Then $a,b\in S_{0}$ and $\mathbf{m}|\mathbf{n}$, that is, $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ must be associative.
As mentioned, having such a complete understanding of all homomorphisms between Petit rings has significant impact to the classification of various associated error-correcting codes. In addition to the conjecture above, we identify that it is possible to extend the results of our paper to the case of noncommutative rings. It would also be interesting to see how our methods can be applied when we look at homomorphisms between general Petit rings.
7. Compliance with ethical standards
Funding sources
This paper was written during the second author’s stay as a Simons Professor in Residence at the University of Ottawa. She gratefully acknowledges the support of CRM and the Simons Foundation, and thanks the Department of Mathematics and Statistics for its hospitality and its congenial and inspiring atmosphere.
Disclosure of potential conflicts of interest
The authors have no relevant financial or non-financial interests to disclose.
Data Availability Statement
No data were generated or used for this research.